# 3D Unsupervised Learning by Distilling 2D Open-Vocabulary Segmentation Models for Autonomous Driving

### v1

@@ 5 Conclusion
We propose UOV, a versatile two-stage unsupervised framework that serves for both 3D pre-training and annotation-free semantic segmentation, achieving state-of-the-art performance across multiple experiments. The key to UOV is to leverage the high-quality knowledge of 2D open-vocabulary segmentation models. Moreover, We propose Tri-Modal contrastive Pre-training (TMP) and Approximate Flat Interaction (AFI) for the first time.
@@
Appendix A Approximate Flat Interaction (AFI)
@@
Firstly, we remove the predictions in the label-confused area. Then, we utilize pseudo-labels ${L}_{\mathcal{P}}^{\text{pseudo}}$ generated in Sec. 3.2 to randomly cover model-predicted labels (note that the coverage step is optional). The probability $P(d_{p})$ for random covering point $p$ is defined as follows:
@@
where $\phi$ is the golden ratio. Normal vectors $FL$ is shown in Fig. 4a.
@@
$C_{n}={argmax}_{i\in 0,\ldots,M-1}\left(\langle(p_{n}-p_{0}),fl_{i}\rangle\right)=\operatorname*{argmax}_{i\in 0,\ldots,M-1}\left(\frac{(p_{n}-p_{0})\cdot fl_{i}}{|p_{n}-p_{0}|}\right)$ (8)
@@
As shown in Fig. 4b, in downsamplings, assuming the mean direction vectors and correlations for two interacting point clouds $\mathcal{P}_{a}$, $\mathcal{P}_{b}$ are ($D_{a}$, $L_{a}$) and ($D_{b}$, $L_{b}$), respectively. Assuming the direction from the center $p_{a0}$ of point cloud $\mathcal{P}_{a}$ to the center $p_{b0}$ of point cloud $\mathcal{P}_{b}$ is the positive direction, we calculate the cosine similarity $S_{\cdot,\cdot}^{a\rightarrow b}$ between the vector connecting two central points and the mean direction vectors $d_{i}\in D_{a}$, $d_{j}\in D_{b}$:
@@
Through Eq. 10,11,12,13,14,15,16, we use the mean direction vectors $\{D_{n}|n=1,\ldots,N\}$, correlations $\{L_{n}|n=1,\ldots,N\}$ and features $\{f_{p_{n}}|n=1,\ldots,N\}$ generated before each downsampling step to assist in generating the new correlation $l_{n_{0},k}$ for each point $p_{k}$ around $p_{n_{0}}$. Then we use Eq. 9 to obtain $L_{n_{0}}^{\prime}=\{l_{i}^{n_{0}}|i=0,\ldots,M-1\}$ from $\{l_{n_{0},k}|p_{k}\in\mathcal{P}_{K_{n_{0}}}\}$ and $F_{n_{0}}^{\prime}=\{f_{i}^{n_{0}}|i=0,\ldots,M-1\}$ from $\{f_{p_{k}}\}$. Afterwords $L_{n_{0}}^{\prime}$ acts as weights for the directional features $F_{n_{0}}^{\prime}$ and contributes to the generation of central point features (Eq. 17). Throughout the entire process, we bind the correlation between two points to 1) whether they are relevant in the same direction (corresponding to $l_{a,b}^{c,*}$ in Eq. 13), and 2) how close the relevance is between the correlated directions (corresponding to $l_{a,b}^{d,*}$ in Eq. 14). Through four rounds of downsampling, point-to-point interactions construct an approximate flat network, thus tightly controlling feature interactions among points (Fig. 4c).
Neural Network Architecture:
@@
$f_{p_{n_{0}}}^{\prime}$ $=\text{softmax}\left(\sum_{i=0}^{M-1}\left(l_{i}^{n_{0}}\times f_{i}^{n_{0}}+0.1\times f_{p_{n_{0}}}^{*}+1\times 10^{-8}\times\text{Maxpooling}_{k\in K_{n_{0}}}\left(f_{p_{k}}^{*}\right)\right)\right),$ (17)
@@
$f_{p_{n_{0}}}^{\prime\prime}=\text{softmax}\left(\sum_{k\in K_{n_{0}}^{\prime}}\operatorname{softmax}_{k\in K_{n_{0}}^{\prime}}\left(\sum_{\left|S_{n_{0},i}^{{n_{0}}\rightarrow k}\right|>\gamma}^{{i\in 0,\ldots,M-1}}l_{i}^{n_{0},k}\right)f_{p_{k}}^{**}+f_{p_{n_{0}}}^{\prime}\right),$ (18)
@@
Appendix B Ablation Study
Table 3: Ablation study on different teacher models of UOV’s annotation-free semantic segmentation(mIoU) on nuScenes [8].
@@ Appendix B Ablation Study
MaskCLIP(ECCV’22) [15] 25.49 - 30.35 27.76 - -
FC-CLIP(NeurIPS’23) [16] 39.00 41.70 41.89 42.44 43.28 43.64
CAT-Seg(arXiv’23) [18] 38.45 40.92 41.79 42.50 42.83 43.64
SAN(CVPR’23) [17] 44.16 47.42 46.26 46.92 47.73 47.89
Table 4: Open-vocabulary semantic segmentation performance on COCO [52] and pseudo-label accuracy(Acc) of different 2D open-vocabulary segmentation models.
@@
MaskCLIP(ECCV’22) [15] 8.2 10.0 23.7 45.9 Semantic 55.00 50.18
FC-CLIP(NeurIPS’23) [16] 14.8 18.2 34.1 58.4 panoptic 45.95 78.96
CAT-Seg(arXiv’23) [18] 10.8 20.4 31.5 62.0 Semantic 55.00 77.05
SAN(CVPR’23) [17] 13.7 17.1 33.3 60.2 Semantic 55.00 80.38
For the ablation study on different setups of UOV (Tab. 3), we utilized MaskCLIP [15] as the control group, and FC-CLIP [16], CAT-Seg [18], and SAN [17] as the experimental groups. Several sets of experiments were conducted while controlling for irrelevant variables. Meanwhile, for a deeper understanding of the impact of certain ablation targets, we conducted separate experiments employing CAT-Seg as the teacher model, as illustrated in the Tab. 5.
Performance Gap in Transferring 2D Open-Vocabulary Segmentation Models to 3D Tasks As shown in Tab. 3, when conducting ablation experiments with different experimental groups, SAN consistently outperforms FC-CLIP and CAT-Seg under any configuration, achieving a 44.16% mIoU in the UOV-baseline, which is 5.16% mIoU and 5.71% mIoU higher than the latter two. We record the accuracy (Acc) of different 2D open-vocabulary segmentation models in generating point cloud pseudo-labels in Tab. 4, while also providing their performance on open-vocabulary segmentation in COCO [52]. We believe that the reasons different teacher models affect student models performance are 1) the ability to comprehend unknown classes, which can affect the accuracy of pseudo-labels, as the number of unknown classes in the class dictionary $\mathcal{C}$ is far greater than that of known classes; Meanwhile, 2) the degree of alignment between image and text features, with SAN exhibiting the largest disparity, making the fitting of the objective loss Eq. 1,2 more challenging during TMP. It is worth noting that models trained with FC-CLIP and CAT-Seg, even after TMP and AFI, do not outperform the baseline of SAN on annotation-free segmentation. This indicates that A Great Teacher Model is All You Need.
The Role of Class Dictionary The class dictionary serves to bridge the datasets’ stuff-classes with open vocabularies. Directly inputting semantic classes of the dataset as prompts into 2D open-vocabulary segmentation models is inefficient, as open-vocabulary models are better suited to understand prompts with smaller semantic granularity. As observed in Tab. 5, not using a class dictionary resulted in a -22.85% mIoU decrease compared to using one.
Table 5: Ablation study of different targets Based on CAT-Seg [18]. $M_{I}$ and $F_{M}$ respectively denote whether employing segmentation results or mask features from CAT-Seg. If $F_{M}$ is not utilized, ResNet50 [46] is integrated as the image backbone. $\mathcal{L}_{T-P}$ indicates the adoption of text-superpoint contrastive loss. $\mathcal{C}$ represents class dictionary, while $+2D$ indicates the incorporation of image inference in AFI.
@@
Ablation Targets in TMP In Tab. 5, employing CAT-Seg [18] as the teacher model, we investigate the efficacy of TMP through two ablation experiments: 1) replacing image features with the outputs of self-supervised ResNet50 [46] backbone using MoCov2 [53, 54], and 2) excluding text-point contrastive loss $\mathcal{L}_{T-P}$. Results from fine-tuning with 1% data and annotation-free point cloud segmentation experiments demonstrate the superiority of TMP over both 1) and 2). Moreover, we affirm the positive impact of TMP on the task of annotation-free point cloud segmentation. As we can see from Tab. 3, incorporating TMP into baselines of FC-CLIP, CAT-Seg, and SAN improved their respective scores by 2.7% mIoU, 2.47% mIoU, and 3.26% mIoU, compared to pre-training with SLidR [39].
@@
Effect of AFI AFI leverages 3D spatial information to enhance the segmentation of unannotated point clouds. Employing AFI on models of FC-CLIP, CAT-Seg, and SAN reached 41.89%, 41.79%, and 46.26% mIoU, a boost of 2.89%, 3.34%, and 2.1% mIoU over their baseline (Tab. 3). It is noteworthy that the combined usage of TMP and AFI results in a greater improvement compared to using either one alone. However, this improvement is not linearly additive.
@@
The Role of Pseudo-Label Guided Knowledge Distillation in Fine-tuning Tasks It is important to note that if we treat UOV as a pre-training method, it does not utilize the validation set in either stage. The two-stage UOV outperforms the one-stage UOV-TMP for downstream tasks on nuScenes, as shown in Tab. 2. The superiority arises from the second-stage annotation-free training and the downstream tasks employing the same loss function, which is particularly effective during linear probing. However, annotation-free training weakens the model’s generalization ability, resulting in inferior performance when transferred to other datasets, e.g., downstream tasks on SemanticKITTI (pre-trained on nuScenes) compared to using only TMP.
@@ Appendix C Liminations
Appendix D Attached Diagrams

### v3

@@ Conclusion
We propose AFOV, a versatile two-stage annotation-free framework that serves for both 3D pre-training and annotation-free semantic segmentation, achieving state-of-the-art performance across multiple experiments. The key to AFOV is to leverage the high-quality knowledge of 2D open-vocabulary segmentation models. Moreover, We propose Tri-Modal contrastive Pre-training (TMP) and Approximate Flat Interaction (AFI) for the first time.
@@
Acknowledgments
This work is supported by the Beijing Natural Science Foundation (L245025) and the Joint Development of Multi-modal Parallel LiDARs with Waytous Inc.
Appendix A A: Details of AFI
@@
Firstly, we remove the predictions in the label-confused area. Then, we utilize pseudo-labels ${L}_{\mathcal{P}}^{\text{pseudo}}$ generated in Baseline of AFOV to randomly cover model-predicted labels (note that the coverage step is optional). The probability $P(d_{p})$ for random covering point $p$ is defined as follows:
@@
where $\phi$ is the golden ratio. Normal vectors $FL$ is shown in Fig. 6 (a).
@@
$C_{n}$ $=\operatorname*{argmax}_{i=0,\ldots,M-1}\left(\langle(p_{n}-p_{0}),fl_{i}\rangle\right)$ (8)
$=\operatorname*{argmax}_{i=0,\ldots,M-1}\left(\frac{(p_{n}-p_{0})\cdot fl_{i}}{|p_{n}-p_{0}|}\right)$
@@
As shown in Fig. 6 (b), in downsamplings, assuming the mean direction vectors and correlations for two interacting point clouds $\mathcal{P}_{a}$, $\mathcal{P}_{b}$ are ($D_{a}$, $L_{a}$) and ($D_{b}$, $L_{b}$), respectively. Assuming the direction from the center $p_{a0}$ of point cloud $\mathcal{P}_{a}$ to the center $p_{b0}$ of point cloud $\mathcal{P}_{b}$ is the positive direction, we calculate the cosine similarity $S_{\cdot,\cdot}^{a\rightarrow b}$ between the vector connecting two central points and the mean direction vectors $d_{i}\in D_{a}$, $d_{j}\in D_{b}$:
@@
Through Eq. 10,11,12,13,14,15,16, we use the mean direction vectors $\{D_{n}|n=1,\ldots,N\}$, correlations $\{L_{n}|n=1,\ldots,N\}$ and features $\{f_{p_{n}}|n=1,\ldots,N\}$ generated before each downsampling step to assist in generating the new correlation $l_{n_{0},k}$ for each point $p_{k}$ around $p_{n_{0}}$. Then we use Eq. 9 to obtain $L_{n_{0}}^{\prime}=\{l_{i}^{n_{0}}|i=0,\ldots,M-1\}$ from $\{l_{n_{0},k}|p_{k}\in\mathcal{P}_{K_{n_{0}}}\}$ and $F_{n_{0}}^{\prime}=\{f_{i}^{n_{0}}|i=0,\ldots,M-1\}$ from $\{f_{p_{k}}\}$. Afterwords $L_{n_{0}}^{\prime}$ acts as weights for the directional features $F_{n_{0}}^{\prime}$ and contributes to the generation of central point features (Eq. 17). Throughout the entire process, we bind the correlation between two points to 1) whether they are relevant in the same direction (corresponding to $l_{a,b}^{c,*}$ in Eq. 13), and 2) how close the relevance is between the correlated directions (corresponding to $l_{a,b}^{d,*}$ in Eq. 14). Through four rounds of downsampling, point-to-point interactions construct an approximate flat network, thus tightly controlling feature interactions among points (Fig. 6 (c)).
Neural Network Architecture
@@
$f_{p_{n_{0}}}^{\prime}=\operatorname{softmax}$ $\left(\sum_{i=0}^{M-1}\left(l_{i}^{n_{0}}\times f_{i}^{n_{0}}+0.1\times f_{p_{n_{0}}}^{*}\right.\right.$ (17)
$\left.+1\times 10^{-8}\times\operatorname*{Maxpooling}_{k\in K_{n_{0}}}\left(f_{p_{k}}^{*}\right)\Big{)}\right),$
@@
$f_{p_{n_{0}}}^{\prime\prime}=\operatorname{softmax}\left(\sum_{k\in K_{n_{0}}^{\prime}}\operatorname*{softmax}_{k\in K_{n_{0}}^{\prime}}\left(\sum_{\left|S_{n_{0},i}^{{n_{0}}\rightarrow k}\right|>\gamma}^{{i\in 0,\ldots,M-1}}l_{i}^{n_{0},k}\right)f_{p_{k}}^{**}\right.$ (18)
$+f_{p_{n_{0}}}^{\prime}\Bigg{)},$
@@
Table 3: Ablation study on different teacher models of AFOV’s annotation-free semantic segmentation(mIoU) on nuScenes (Caesar et al. 2020).
@@
MaskCLIP(ECCV’22) (Zhou, Loy, and Dai 2022) 25.49 - 30.35 27.76 - -
FC-CLIP(NeurIPS’23) (Yu et al. 2023) 39.00 41.70 41.89 42.44 43.28 43.64
CAT-Seg(CVPR’24) (Cho et al. 2023) 38.45 40.92 41.79 42.50 42.83 43.64
SAN(CVPR’23) (Xu et al. 2023) 44.16 47.42 46.26 46.92 47.73 47.89
Table 4: Open-vocabulary semantic segmentation performance on COCO (Lin et al. 2014) and pseudo-label accuracy(Acc) of different 2D open-vocabulary segmentation models.
@@
MaskCLIP(ECCV’22) (Zhou, Loy, and Dai 2022) 8.2 10.0 23.7 45.9 Semantic 55.00 50.18
FC-CLIP(NeurIPS’23) (Yu et al. 2023) 14.8 18.2 34.1 58.4 panoptic 45.95 78.96
CAT-Seg(CVPR’24) (Cho et al. 2023) 10.8 20.4 31.5 62.0 Semantic 55.00 77.05
SAN(CVPR’23) (Xu et al. 2023) 13.7 17.1 33.3 60.2 Semantic 55.00 80.38
Table 5: Ablation study of different targets based on CAT-Seg (Cho et al. 2023). $M_{I}$ and $F_{M}$ respectively denote whether employing segmentation results or mask features from CAT-Seg. If $F_{M}$ is not utilized, ResNet50 (He et al. 2016) is integrated as the image backbone. $\mathcal{L}_{T-P}$ indicates the adoption of text-superpoint contrastive loss. $SP$ means whether to use superpixel and superpoint. $\mathcal{C}$ represents class dictionary, while $+2D$ indicates the incorporation of image inference in AFI.
@@
Table 6: Ablation Study on 2D Feature Pre-generation. Experiments were conducted based on whether the mask features were pre-generated during CAT-Seg pre-training (if the features were not pre-generated, the 2D backbone network in TMP exists and all layers except the feature output layer are frozen). SLidR was used as the control group.
Pre-train Method Annotation-free 1% Fine-tuning
Pre-training
Time (h/epoch)
SLidR 38.45 38.30 2.82
TMP(w/ 2D backbone) 40.12 46.80 3.44
TMP(w/o 2D backbone) 40.92 46.61 2.10
Appendix B B: Ablation Study
For the ablation study on different setups of AFOV (Tab. 3), we utilized MaskCLIP (Zhou, Loy, and Dai 2022) as the control group, and FC-CLIP (Yu et al. 2023), CAT-Seg (Cho et al. 2023), and SAN (Xu et al. 2023) as the experimental groups. Several sets of experiments were conducted while controlling for irrelevant variables. Meanwhile, for a deeper understanding of the impact of certain ablation targets, we conducted separate experiments employing CAT-Seg as the teacher model, as illustrated in the Tab. 5 and Tab. 6.
Performance Gap in Transferring 2D Open-Vocabulary Segmentation Models to 3D Tasks
As shown in Tab. 3, when conducting ablation experiments with different experimental groups, SAN consistently outperforms FC-CLIP and CAT-Seg under any configuration, achieving a 44.16% mIoU in the AFOV-baseline, which is 5.16% mIoU and 5.71% mIoU higher than the latter two. We record the accuracy (Acc) of different 2D open-vocabulary segmentation models in generating point cloud pseudo-labels in Tab. 4, while also providing their performance on open-vocabulary segmentation in COCO (Lin et al. 2014). We believe that the reasons different teacher models affect student models performance are 1) the ability to comprehend unknown classes, which can affect the accuracy of pseudo-labels, as the number of unknown classes in the class dictionary $\mathcal{C}$ is far greater than that of known classes; Meanwhile, 2) the degree of alignment between image and text features, with SAN exhibiting the largest disparity, making the fitting of the objective loss Eq. 1,2 more challenging during TMP. It is worth noting that models trained with FC-CLIP and CAT-Seg, even after TMP and AFI, do not outperform the baseline of SAN on annotation-free segmentation. This indicates that A Great Teacher Model is All You Need.
The Role of Class Dictionary
The class dictionary serves to bridge the datasets’ stuff-classes with open vocabularies. Directly inputting semantic classes of the dataset as prompts into 2D open-vocabulary segmentation models is inefficient, as open-vocabulary models are better suited to understand prompts with smaller semantic granularity. As observed in Tab. 5, not using a class dictionary resulted in a -6.14% mIoU decrease compared to using one.
Ablation Targets in TMP
@@ Appendix B B: Ablation Study
In Tab. 5, employing CAT-Seg (Cho et al. 2023) as the teacher model, we investigate the efficacy of TMP through two ablation experiments: 1) replacing image features with the outputs of self-supervised ResNet50 (He et al. 2016) backbone using MoCov2 (Chen et al. 2020; He et al. 2020), and 2) excluding text-point contrastive loss $\mathcal{L}_{T-P}$. Results from fine-tuning with 1% data and annotation-free point cloud segmentation experiments demonstrate the superiority of TMP over both 1) and 2). Moreover, we affirm the positive impact of TMP on the task of annotation-free point cloud segmentation. As we can see from Tab. 3, incorporating TMP into baselines of FC-CLIP, CAT-Seg, and SAN improved their respective scores by 2.7% mIoU, 2.47% mIoU, and 3.26% mIoU, compared to pre-training with SLidR (Sautier et al. 2022).
The Role of Superpoint-superpixel in TMP
Superpoint and superpixel can avoid self-conflict among point groups (such as $k$-NN groups) with the same semantics. In Tab. 5, $SP$ represents whether superpixel-superpoint is used; otherwise, we use $k$-NN to divide up point cloud. The experimental results show that using superpixel-superpoint leads to performance improvements of +1.58% mIoU in the annotation-free segmentation and +4.74% mIoU in the 1% fine-tuning task.
Feature Pre-generation
In the experiment of adding a 2D backbone in Tab. 6, similar to SLidR (Sautier et al. 2022), we only activated the output layer of the added 2D backbone (with other layers frozen). The results show that TMP (w/o 2D backbone) has a 25% and 39% improvement in runtime over SLidR and TMP (w/ 2D backbone). Moreover, adding a 2D network does not provide performance advantages to TMP. At the same time, the storage burden of pre-generated features is negligible (¡7GB). After considering all factors, we retained the step of pre-generating 2D features.
Effect of AFI
@@
AFI leverages 3D spatial information to enhance the segmentation of unannotated point clouds. Employing AFI on models of FC-CLIP, CAT-Seg, and SAN reached 41.89%, 41.79%, and 46.26% mIoU, a boost of 2.89%, 3.34%, and 2.1% mIoU over their baseline (Tab. 3). It is noteworthy that the combined usage of TMP and AFI results in a greater improvement compared to using either one alone. However, this improvement is not linearly additive.
The Role of Pseudo-Label Guided Knowledge Distillation in Fine-tuning Tasks
@@
It is important to note that if we treat AFOV as a pre-training method, it does not utilize the validation set in either stage. The two-stage AFOV outperforms the one-stage AFOV-TMP for downstream tasks on nuScenes, as shown in Tab. 2. The superiority arises from the second-stage annotation-free training and the downstream tasks employing the same loss function. Essentially, this is over-fitting on the nuScenes data, which is particularly effective during linear probing. However, annotation-free training weakens the model’s generalization ability, resulting in inferior performance when transferred to other datasets, e.g., downstream tasks on SemanticKITTI (pre-trained on nuScenes) compared to using only TMP.
@@ Appendix C C: Liminations
Appendix D D: Attached Diagrams